### v2

# ReCalKV: Low-Rank KV Cache Compression via Head Reordering and Offline Calibration

Abstract
Large language models (LLMs) have achieved remarkable performance, yet their capability on long-context reasoning is often constrained by the excessive memory required to store the Key-Value (KV) cache. This makes KV cache compression an essential step toward enabling efficient long-context reasoning. Recent methods have explored reducing the hidden dimensions of the KV cache, but many introduce additional computation through projection layers or suffer from significant performance degradation under high compression ratios. To address these challenges, we propose ReCalKV, a post-training KV cache compression method that reduces the hidden dimensions of the KV cache. We develop distinct compression strategies for Keys and Values based on their different roles and varying importance in the attention mechanism. For Keys, we propose Head-wise Similarity–aware Reordering (HSR), which clusters similar heads and applies grouped SVD to the key projection matrix, reducing additional computation while preserving accuracy. For Values, we propose Offline Calibration and Matrix Fusion (OCMF) to preserve accuracy without extra computational overhead. Experiments show that ReCalKV outperforms existing low-rank compression methods, achieving high compression ratios with minimal performance loss. The code and models will be available at: https://github.com/XIANGLONGYAN/ReCalKV.
1 Introduction
Large language models (LLMs) vaswani2017attention ; llama1 ; llama3 have demonstrated outstanding performance across a wide range of tasks. To accelerate inference, modern LLMs cache intermediate Key-Value (KV) states, avoiding redundant computation during autoregressive decoding. However, as the input context length increases, the KV cache grows rapidly, leading to substantial memory overhead and bandwidth pressure. In practice, the KV cache often becomes the primary bottleneck for long-context inference. Consequently, compressing the KV cache is crucial for efficient and scalable deployment of LLMs.
To reduce the size of the KV cache, recent works explore compression along multiple axes. Multi-query attention mqa and grouped-query attention ainslie2023gqa reduce the number of heads by sharing keys and values. Quantization methods atom lower KV cache’s precision, with some kivi ; kvquant pushing KV representations down to 2 bits. Others h2o ; snapkv ; xiao2023efficient ; dong2024get reduce the number of cached tokens by selecting only important ones, often based on attention scores. A few methods chang2025xkv ; liu2024minicache further compress across layers by reusing KV states. These approaches highlight the multi-dimensional nature of KV cache compression, targeting heads, precision, sequence length, and depth.
Another line of work chang2024palu ; liu2024deepseek ; lin2024matryoshkakv explores KV cache compression from a different angle—by reducing the dimensionality of the hidden vector space of Keys and Values themselves. For example, MLA liu2024deepseek reduces memory via low-rank representations but requires training the model from scratch. Other approaches, such as EigenAttention saxena2024eigen and MastryohakaKV lin2024matryoshkakv , compress KV entries via projection into lower-dimensional subspaces. Although effective, their projection and reconstruction introduce extra decoding overhead, limiting applicability in latency-sensitive scenarios. Palu chang2024palu and LoRC zhang2024lorc address this issue by directly applying Singular Value Decomposition (SVD) to the KV projection layers, effectively compressing the hidden dimensions of the KV cache. This approach substantially reduces the overhead of runtime projection and reconstruction. However, both methods overlook the inherent asymmetry between Keys and Values in the attention mechanism, and their performance degrades notably under high KV cache compression ratios.
To better explore KV cache compression along the hidden dimension, we conduct detailed analyses of the roles of Keys and Values in the attention mechanism. Our analyses reveal that: (i) Most modern LLMs llama1 ; llama2 ; llama3 use positional encoding—typically RoPE rope_roformer —which is applied to Keys. As a result, low-rank compressed Keys must be fully reconstructed during inference to enable positional encoding, which introduces additional computational overhead. This makes it essential to consider both accuracy and computational cost when compressing Key cache. (ii) We measure the Fisher information of the Key and Value projection layers. Our Fisher information analysis reveals that the Value projection matrices carry significantly higher importance than their Key counterparts, highlighting the crucial role of Value representations in the overall model behavior. Therefore, when compressing the Value cache, it is crucial to minimize accuracy degradation.
Based on our analyses, we develop distinct compression strategies for Keys and Values based on their different roles and varying importance in the attention mechanism. For Keys, we propose Head-wise Similarity–aware Reordering (HSR). It first reorders attention heads based on their representation similarity, then groups similar heads together, and applies grouped SVD within each group. This reduces the Key cache size with low reconstruction overhead. Grouping similar heads helps the SVD better capture shared subspace structures, lowering approximation error and preserving accuracy. For Values, we propose Offline Calibration and Matrix Fusion (OCMF). We first apply SVD to the Value projection matrix, then calibrate the decomposed components using a small calibration dataset. Finally, we fuse the right factor of the SVD decomposition with the subsequent output projection matrix, eliminating the need for explicit reconstruction during inference.
Extensive experiments demonstrate that ReCalKV consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing low-rank compression methods under various evaluation settings. For example, on the LLaMA-2-7B model llama2 evaluated on six zero-shot QA datasets, ReCalKV achieves an average accuracy of 63.64% under a 50% KV cache compression ratio, compared to 64.99% for the full-precision model—corresponding to only a 2% relative drop. Notably, since ReCalKV is orthogonal to quantization techniques, it can be seamlessly integrated with them to achieve even higher overall compression ratios.
Our key contributions can be summarized as follows:
  • We propose ReCalKV, a novel post-training KV cache compression framework that reduces memory overhead in long-context inference without requiring model retraining.
  • We propose the Head-wise Similarity–aware Reordering (HSR) strategy for Key compression, which effectively reduces the Key cache size with limited reconstruction overhead.
  • We propose the Offline Calibration and Matrix Fusion (OCMF) strategy for Value compression, preserving accuracy without added computation.
  • Extensive experiments demonstrate that ReCalKV consistently outperforms prior low-rank compression approaches. Furthermore, ReCalKV can be seamlessly combined with quantization techniques to achieve even higher compression ratios.
2 Related work
SVD-Based LLM Compression. Singular Value Decomposition (SVD) has been widely adopted for compressing LLMs by approximating weight matrices with low-rank factors. However, standard SVD NaiveSVD has been observed to cause notable performance degradation in practice. To address this, FWSVDfwsvd incorporates Fisher information to guide decomposition, while ASVDyuan2023asvd rescales weights to mitigate the impact of activation outliers. SVD-LLMsvdllm establishes a direct connection between singular values and compression loss, employing data whitening and selective truncation to minimize degradation. AdaSVDli2025adasvd further improves compression by introducing adaptive rank allocation and compensation mechanisms. Unlike prior methods that target model weights, we apply SVD to the key and value projections to compress the KV cache directly.
KV Cache Compression. To support long-context inference, various methods have been proposed to compress the KV cache. Quantization-based approaches are widely adopted: Atomatom performs per-token quantization, while WKVQuantwkvquant uses a two-level scheme for better accuracy. KIVIkivi and KVQuantkvquant combine per-token and per-channel quantization, with KVQuant further leveraging non-uniform and sparse techniques to handle outliers. In parallel, token eviction methods h2o ; snapkv ; xiao2023efficient ; dong2024get reduce memory by discarding less relevant tokens or retrieving only subsets during decoding. Beyond these strategies, several methods aim to reduce the hidden dimension of KV representations. DeepSeekV2liu2024deepseek uses MLA for built-in dimension reduction, but requires training from scratch. MatryoshkaKVlin2024matryoshkakv and Eigen-Attentionsaxena2024eigen project the KV cache into a low-rank space via additional projection matrices, at the cost of increased computation. HeadKV fu2024not uses SVD to reduce the number of KV heads, leading to improved efficiency. LoRCzhang2024lorc and Paluchang2024palu directly apply SVD to the KV projection matrices, reducing dimensionality with minimal architectural changes. Our method is a post-training low-rank KV compression approach and is orthogonal to quantization and token eviction, enabling easy integration for further compression gains.
Figure 1: Overview of the ReCalKV framework. The method consists of three key components: Head-wise Similarity-aware Reordering (HSR) for compressing Keys via grouped SVD, and Offline Calibration and Matrix Fusion (OCMF) for compressing Values without additional runtime overhead.
3 Methodology
3.1 Preliminary
Singular Value Decomposition. Singular Value Decomposition svd is a classical matrix factorization technique widely used for low-rank approximation. Given a matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD factorizes it into three components: $\mathbf{W}=\mathbf{U}{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, and $\boldsymbol{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix with non-negative singular values. To obtain a low-rank approximation of a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, we apply singular value decomposition (SVD) and retain only the top $r$ singular values and their associated singular vectors. This results in an approximate factorization:
$\mathbf{W}\approx\mathbf{L}\mathbf{R},\quad\text{where}\quad\mathbf{L}=\mathbf{U}_{r}\boldsymbol{\Sigma}_{r}^{1/2},\quad\mathbf{R}=\boldsymbol{\Sigma}_{r}^{1/2}\mathbf{V}_{r}^{\top}.$ (1)
Here, $\mathbf{U}_{r}\in\mathbb{R}^{m\times r}$ and $\mathbf{V}_{r}\in\mathbb{R}^{n\times r}$ are the top-$r$ singular vectors of $\mathbf{W}$, and $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$ contains the corresponding singular values. This yields two smaller matrices, $\mathbf{L}$ and $\mathbf{R}$, from the low-rank decomposition of $\mathbf{W}$. Given an input $\mathbf{x}\in\mathbb{R}^{1\times m}$, we compute the intermediate representation $\mathbf{z}=\mathbf{x}\mathbf{L}$ and store $\mathbf{z}\in\mathbb{R}^{1\times r}$ in the KV cache instead of the full output. During attention, the output is approximated by reconstructing $\mathbf{x}\mathbf{W}\approx\mathbf{z}\mathbf{R}$,. This approach reduces the KV cache size with a compression ratio of $r/n$, while maintaining a close approximation of the original computation.
Centered Kernel Alignment Similarity. Centered Kernel Alignment (CKA) kornblith2019similarity is a widely used metric for quantifying the similarity between two sets of representations. Given two matrices $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, CKA is computed by first forming their Gram (kernel) matrices $\mathbf{G_{X}}=\mathbf{X}\mathbf{X}^{\top}$ and $\mathbf{G_{Y}}=\mathbf{Y}\mathbf{Y}^{\top}$. These kernel matrices are then centered as follows:
$\mathbf{\tilde{G}_{X}}=\mathbf{H}\mathbf{G_{X}}\mathbf{H},\quad\mathbf{\tilde{G_{Y}}}=\mathbf{H}\mathbf{G_{Y}}\mathbf{H},$ (2)
where $\mathbf{H}=\mathbf{I}_{n}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$ is the centering matrix. The final CKA similarity is defined as:
$\text{CKA}(\mathbf{X},\mathbf{Y})=\frac{\text{HSIC}(\mathbf{X},\mathbf{Y})}{\sqrt{\text{HSIC}(\mathbf{X},\mathbf{X})\cdot\text{HSIC}(\mathbf{Y},\mathbf{Y})}},$ (3)
where $\text{HSIC}(\mathbf{X},\mathbf{Y})=\text{Tr}(\mathbf{\tilde{G_{X}}}\mathbf{\tilde{G_{Y}}})$ denotes the Hilbert-Schmidt Independence Criterion (HSIC). CKA ranges from 0 to 1, with higher values indicating greater similarity between the two matrics.
3.2 Head-wise Similarity-aware Reordering
Figure 2: CKA similarity matrices before and after head reordering.
Group-head Low-rank Decomposition. Following the grouped decomposition strategy proposed in Palu chang2024palu , we organize multiple attention heads into groups prior to performing SVD. Given a Key projection matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, where $n=h\cdot d_{h}$ corresponds to $h$ attention heads each of hidden dimension $d_{h}$, we divide $\mathbf{W}$ column-wise into $h$ submatrices, each representing a single head. Then we group every $s$ heads into one group, resulting in $g=h/s$ groups in total. For each group $j$, we construct a concatenated projection matrix $\mathbf{W}_{g_{j}}=\bigl{[}\mathbf{W}_{j,1},\dots,\mathbf{W}_{j,s}\bigr{]}$, where each $\mathbf{W}_{j,k}\in\mathbb{R}^{m\times d_{h}}$ is the projection matrix of the $k$-th head in group $j$, and thus $\mathbf{W}_{g_{j}}\in\mathbb{R}^{m\times(s\cdot d_{h})}$. Instead of applying SVD to the entire projection matrix at once, we apply low-rank approximation to the grouped matrix: $\mathbf{W}_{g_{j}}\approx\mathbf{L}_{g_{j}}\mathbf{R}_{g_{j}}$, where $\mathbf{L}_{g_{j}}\in\mathbb{R}^{d\times r_{g}}$ and $\mathbf{R}_{g_{j}}\in\mathbb{R}^{r_{g}\times(d_{h}\cdot s)}$. During inference, the latent representation shared across all heads in the group is computed as: $\mathbf{z}_{g_{j}}=\mathbf{x}\mathbf{L}_{g_{j}}$, and the projected outputs for individual heads are reconstructed via:
$[\mathbf{y}_{j,1},\dots,\mathbf{y}_{j,s}]=\mathbf{z}_{g_{j}}\mathbf{R}_{g_{j}}.$ (4)
This grouped strategy provides a good trade-off between reconstruction overhead and approximation fidelity, enabling efficient compression with minimal performance impact.
Figure 3: Key decoding with HSR. Similar heads are reordered and grouped before SVD, enabling more accurate reconstruction.
CKA-based Head Similarity. A key question in grouped SVD is how to assign attention heads into groups so as to minimize the reconstruction error. Empirically, we observe that grouping heads with similar left singular subspaces results in lower approximation error, as these heads tend to share more common representational components and thus benefit from joint compression. To quantify head similarity, we adopt centered kernel alignment (CKA) kornblith2019similarity , a robust and widely used metric for comparing representation subspaces. Specifically, we compute the pairwise CKA similarity between all attention heads, yielding a symmetric similarity matrix $\mathbf{S}\in\mathbb{R}^{h\times h}$, where $h$ is the total number of heads:
$\mathbf{S}_{i,j}=\text{CKA}(\mathbf{H}_{i},\mathbf{H}_{j}),\quad\forall i,j\in{1,\dots,h}.$ (5)
Head Reordering. Based on the similarity matrix $\mathbf{S}$, we perform head grouping by prioritizing pairs with high mutual similarity. Specifically, we adopt a greedy strategy that iteratively selects the head pair with the highest CKA similarity and assigns them to the same group, subject to a fixed group size constraint (e.g., 4 heads per group when $h=32$). Remaining unassigned heads are then added to existing groups with available capacity, ensuring that all heads are eventually grouped. This head reordering process encourages heads with similar representational structure to share SVD decompositions, effectively reducing approximation error during grouped compression. As shown in Figure 2, we visualize the CKA similarity matrices before and after head reordering. It can be observed that, after reordering, heads assigned to adjacent positions exhibit higher mutual similarity, indicating that similar heads are more effectively grouped together. By aligning structurally coherent heads within each group, we improve both the compactness and accuracy of the resulting low-rank representation. To ensure decoding equivalence, we apply an inverse reordering to restore the original head order after grouped computation (see Figure 3).
3.3 Offline Calibration and Matrix Fusion
Offline Calibration. For Value cache compression, we directly perform SVD on the full Value projection matrix $\mathbf{W}_{v}\in\mathbb{R}^{m\times n}$, resulting in the decomposed matrices: $\mathbf{W}_{v}\approx\mathbf{L}_{v}\mathbf{R}_{v}$, where $\mathbf{L}_{v}\in\mathbb{R}^{m\times r}$ and $\mathbf{R}_{v}\in\mathbb{R}^{r\times n}$ represent the compressed components of the original Value projection matrix $\mathbf{W}_{v}\in\mathbb{R}^{m\times n}$. We then define the approximation error $\mathcal{E}$ introduced by the SVD decomposition as:
$\mathcal{E}=||\mathbf{L}_{v}\mathbf{R}_{v}\mathbf{X}-\mathbf{W}_{v}\mathbf{X}||_{F}^{2},$ (6)
where $\mathbf{X}$ denotes the calibration dataset. Based on our analysis of Fisher Information, we observe that the Value projection matrix exhibits significantly higher Fisher Information compared to the Key projection matrix. This indicates that the Value projection matrix plays a more critical role in model performance. Therefore, we aim to minimize the approximation error $\mathcal{E}$ introduced during the compression of the Value projection to preserve model accuracy as much as possible. Inspired by recent findings in adaptive decomposition methods such as AdaSVD li2025adasvd , we observe that standard SVD decomposition does not always yield the lowest approximation error. To this end, we perform offline calibration of the decomposed matrices $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ using a small calibration dataset $\mathbf{X}$, aiming to further reduce the compression-induced approximation error. We first calibrate $\mathbf{L}_{v}$ by setting the derivative of the approximation error $\mathcal{E}$ with respect to $\mathbf{L}_{v}$ to zero:
$\frac{\partial\mathcal{E}}{\partial\mathbf{L}_{v}}=0\quad\Rightarrow\mathbf{L}_{v}=\mathbf{WX}\mathbf{X}^{\top}\mathbf{R}_{v}^{\top}(\mathbf{R}_{v}\mathbf{X}\mathbf{X}^{\top}\mathbf{R}_{v})^{-1}.$ (7)
Next, to improve the fidelity of the low-rank approximation, we calibrate the right factor $\mathbf{R}_{v}$ by minimizing the approximation error. Specifically, we compute the gradient of the approximation error $\mathcal{E}$ with respect to $\mathbf{R}_{v}$ and set it to zero, allowing us to solve for a more accurate $\mathbf{R}_{v}$ in closed form:
$\frac{\partial\mathcal{E}}{\partial{\mathbf{R}_{v}}}=0\quad\Rightarrow{\mathbf{R}_{v}}=((\mathbf{L}_{v})^{\top}\mathbf{L}_{v})^{-1}(\mathbf{L}_{v})^{\top}\mathbf{W}.$ (8)
This yields an optimal closed-form solution for $\mathbf{R}_{v}$. Together, the calibrated $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ form a refined low-rank approximation of the original Value projection matrix, effectively reducing compression-induced error without incurring any additional inference-time computation. The derivation of the calibration objective and the resulting closed-form updates for both $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ are provided in the supplementary material, along with a detailed discussion of its implementation considerations.
Algorithm 1 Pseudocode of ReCalKV
1:Inputs: Model $\mathcal{M}$, Calibration Data $\mathcal{X}$, Target Compression Ratio $\mathcal{TR}$
2:Output: $\mathcal{M}^{\prime}$: Model equipped with compressed KV cache
3:procedure ReCalKV($\mathcal{M},\mathcal{X},\mathcal{TR}$)
4:     $\mathcal{F}\leftarrow$ Calculate_Fisher_Info$(\mathcal{M},\mathcal{X})$
5:     $R\leftarrow$ Allocate_Compression_Ratio$(\mathcal{M},\mathcal{TR},\mathcal{F})$
6:     for each Key projection layer $\mathcal{W}_{k}$ in model $\mathcal{M}$ do
7:         $\mathcal{CKA}\leftarrow$ Calculate_CKA_Similarity$(\mathcal{W}_{k})$
8:         $\mathcal{W}_{k}^{\prime}\leftarrow$ Head_Reorder$(\mathcal{W}_{k},\mathcal{CKA})$
9:         $\mathbf{L}_{k}[],\mathbf{R}_{k}[]\leftarrow$ Group_SVD$(\mathcal{W}_{k}^{\prime},\mathcal{R}[\mathcal{W}_{k}])$
10:         $\mathcal{M}^{\prime}\leftarrow$Update_Layer$(\mathbf{L}_{k}[]\&\mathbf{R}_{k}[],\mathcal{W}_{k},\mathcal{M})$
11:     end for
12:     for each Value projection layer $\mathcal{W}_{v}$ in model $\mathcal{M}$ do
13:         $\mathbf{L}_{v},\mathbf{R}_{v}\leftarrow$ SVD$(\mathcal{W}_{v},\mathcal{R}[\mathcal{W}_{v}])$
14:         $\mathbf{L}_{v}^{\prime},\mathbf{R}_{v}^{\prime}\leftarrow$ Offline_Calibration$(\mathcal{W}_{v},\mathbf{L}_{v},\mathbf{R}_{v},\mathcal{X})$
15:         $\widetilde{\mathcal{W}}_{o}\leftarrow$Matrix_Fusion$(\mathbf{R}_{v}^{\prime},\mathcal{W}_{o})$
16:         $\mathcal{M}^{\prime}\leftarrow$Update_Layer$(\mathbf{L}_{v}^{\prime},\mathcal{W}_{v},\mathcal{M})$
17:         $\mathcal{M}^{\prime}\leftarrow$Update_Layer$(\mathcal{W}_{o}^{\prime},\mathcal{W}_{o},\mathcal{M})$
18:     end for
19:     return $\mathcal{M}^{\prime}$
20:end procedure
Matrix Fusion. After performing SVD and offline calibration on the Value projection matrix, we further optimize the inference efficiency by eliminating unnecessary reconstruction steps. Specifically, instead of computing the full Value cache and then applying the output projection matrix $W_{o}$, we fuse $\mathbf{R}_{v}$ into $W_{o}$ to avoid runtime reconstruction. We start from the standard attention output:
$\text{Output}=\text{Attention}(Q,K,V)W_{o}=\text{Attention}(XW_{q},XW_{k},XW_{v})W_{o}.$ (9)
where $X$ is the input sequence, $W_{q}$, $W_{k}$, and $W_{v}$ are the projection matrices for query, key, and value respectively, and $W_{o}$ is the output projection matrix. For compressed Value cache, we store $X^{\prime}=X\mathbf{L}_{v}$ as the low-rank representation. With Value compression, the attention output becomes:
$\text{Output}=\text{Attention}(XW_{q},XW_{k},X\textbf{L}_{v}\textbf{R}_{v})W_{o}.$ (10)
We then define a fused output projection matrix $\widetilde{W}_{o}=\mathbf{R}_{v}W_{o}$, and compute the attention output as:
$\text{Output}=\text{Attention}(XW_{q},XW_{k},X\mathbf{L}_{v})\widetilde{W}_{o}.$ (11)
This design eliminates the need to explicitly reconstruct $X^{\prime}\mathbf{R}_{v}$ during inference. By merging the computation into a single step, matrix fusion reduces both memory usage and computational overhead, streamlining the runtime execution path. Moreover, since the fused matrix $\widetilde{W}_{o}$ can be precomputed entirely offline, it introduces no additional overhead to online inference, making this approach highly suitable for latency-sensitive applications and deployment on resource-constrained hardware.
3.4 ReCalKV Workflow
The overall pipeline of ReCalKV is outlined in Algorithm 1. Given a pre-trained model $\mathcal{M}$, calibration data $\mathcal{X}$, and a target compression ratio $\mathcal{R}$, ReCalKV applies differentiated compression strategies to Key and Value projection layers. We first compute layer-wise Fisher Information scores using the calibration data, following the strategy introduced in Palu chang2024palu , to estimate the relative importance of each layer and guide compression ratio allocation. Based on this, we allocate compression ratios to different layers accordingly. For Key projection layers, we first compute head-wise CKA similarity, then reorder the heads to group the most similar ones together, and finally apply grouped SVD for compression. For Value projection layers, we apply standard SVD followed by offline calibration to adjust the low-rank factors. Then we fuse the right factor of the decomposition with the output projection matrix $W_{o}$ in the attention block. The modified matrices are then written back into the model. The final output is a modified model $\mathcal{M}^{\prime}$ that generates compressed KV caches.
4 Experiments
4.1 Experimental Settings
Models. We evaluate our method on a range of widely adopted LLMs, including multiple generations of the LLaMA family: LLaMA-7B llama1 , LLaMA-2-7B llama2 , and LLaMA-2-13B-Chat llama2 . We also include instruction-tuned variants such as Mistral-7B-Instruct-v0.2 mistral and LongChat-7B-v1.5-32k li2023long to assess performance under extended context settings. These models cover both base and chat/instruction-following variants, ensuring a comprehensive evaluation across different architectures and use cases. Notably, Mistral-7B-Instruct-v0.2 mistral adopt Grouped Query Attention (GQA) ainslie2023gqa , while the others use standard Multi-Head Attention (MHA) vaswani2017attention .
Datasets and Evaluation. We assess the effectiveness of our method using both perplexity and task-specific accuracy. For language modeling evaluation, we report perplexity on WikiText2 wikitext2 , Penn Treebank (PTB)plotz2017benchmarking , and a subset of the C4 corpusraffel2020exploring . To evaluate reasoning and generalization capabilities, we measure zero-shot accuracy on six QA benchmarks: ARC-c, ARC-e clark2018think , Hellaswag zellers2019hellaswag , OBQA openbookqa , PIQA bisk2020piqa , and Winogrande winogrande . In addition, we adopt the LongBench benchmark LongBench to evaluate long-context performance, conducting experiments on eight diverse tasks that comprehensively test the model’s ability to handle extended contexts.
Baselines. We compare our method with Palu chang2024palu , a recent low-rank compression approach for KV cache. Specifically, we adopt its G-LRD variant for evaluation. To ensure a fair comparison, we follow the same group-wise decomposition setting and fix the group size to 4 in all experiments.
Implementation Details. All experiments are conducted using PyTorch paszke2019pytorch and Huggingface Transformers paszke1912imperative on a single NVIDIA A800 GPU with 80GB of memory. Following the setup in SVD-LLM svdllm , we apply a whitening transformation before performing SVD truncation. Specifically, we randomly select 256 samples from the WikiText-2 dataset as calibration data and use them both for whitening in the SVD step and for the offline calibration process in Value compression.
Table 1: Zero-shot performance comparison between ReCalKV and Palu chang2024palu under 50% to 70% compression ratios. Evaluation on three language modeling datasets (measured by perplexity ($\downarrow$)) and six zero-shot QA datasets (measured by both individual and average accuracy ($\uparrow$)).
[…]
LLaMA-7B llama1
0% Original 5.68 41.15 7.34 44.40 76.18 78.67 75.25 44.80 70.01 64.89
50% Palu chang2024palu 6.27 48.39 8.85 41.6 73.46 76.71 72.35 40.53 68.75 62.23
ReCalKV 6.13 43.99 8.36 42.00 73.59 77.20 72.35 41.72 68.35 62.54
60% Palu chang2024palu 7.08 91.45 10.98 36.80 68.80 73.94 67.63 38.14 63.61 58.15
ReCalKV 6.63 56.49 9.44 39.80 71.51 75.95 71.09 40.10 64.17 60.44
70% Palu chang2024palu 8.42 211.33 14.46 34.60 61.06 71.00 61.53 33.70 59.51 53.57
ReCalKV 7.24 71.47 10.53 38.20 68.73 75.19 67.55 39.08 64.01 58.79
LLaMA-2-7B llama2
0% Original 5.47 37.91 7.26 44.20 76.01 78.07 76.35 46.25 69.06 64.99
50% Palu chang2024palu 6.02 40.89 8.72 43.80 73.34 76.17 72.98 42.24 67.32 62.64
ReCalKV 5.83 39.51 8.14 45.00 74.39 76.39 74.71 43.52 67.80 63.64
60% Palu chang2024palu 6.81 51.32 10.69 40.00 68.55 74.10 68.99 38.14 63.38 58.85
ReCalKV 6.21 65.73 8.95 41.40 72.36 76.17 72.73 41.13 68.03 61.97
70% Palu chang2024palu 8.62 83.19 15.01 34.20 59.30 68.82 57.87 31.66 61.01 52.14
ReCalKV 6.75 75.78 10.05 39.80 69.59 74.48 70.37 39.42 65.75 59.90
Mistral-7B-Instruct-v0.2 mistral
0% Original 5.94 32.46 9.72 46.80 83.68 80.41 81.31 55.63 74.35 70.36
50% Palu chang2024palu 6.33 37.38 10.79 44.60 80.89 79.33 79.21 54.27 73.40 68.62
ReCalKV 6.30 38.12 10.73 44.40 81.06 80.20 80.05 54.86 73.56 69.02
60% Palu chang2024palu 7.07 49.45 12.93 42.80 75.30 77.26 74.07 50.00 70.72 65.03
ReCalKV 6.81 47.27 11.98 44.00 77.88 79.27 77.78 52.22 72.30 67.24
70% Palu chang2024palu 8.71 77.51 16.78 38.60 66.48 75.24 66.96 42.49 66.54 59.39
ReCalKV 8.08 70.96 14.98 39.00 71.08 76.82 72.14 46.25 67.72 62.17
LongChat-7B-v1.5-32k li2023long
0% Original 7.61 89.04 10.52 41.40 71.28 76.12 71.84 41.38 68.19 61.70
50% Palu chang2024palu 8.11 120.11 12.08 38.20 68.30 72.52 67.93 38.40 64.96 58.39
ReCalKV 7.89 95.51 11.48 41.80 69.66 73.78 69.61 38.91 65.59 59.89
60% Palu chang2024palu 9.15 168.94 14.42 37.80 64.76 69.70 60.14 34.64 61.09 54.68
ReCalKV 8.14 108.52 12.12 40.00 67.63 71.98 66.92 37.03 63.77 57.89
70% Palu chang2024palu 11.95 172.23 20.87 32.20 54.94 64.74 50.00 28.67 55.56 47.69
ReCalKV 9.01 109.38 13.63 35.20 63.18 68.55 58.84 33.53 59.12 53.07
Table 2: Evaluation results on LongBench LongBench , covering accuracy across 8 tasks and the overall average, comparing ReCalKV and Palu chang2024palu under 50%–70% KV cache compression ratios.
[…]
LLaMA-2-7B llama2
0% Original 9.58 21.22 3.51 66.00 87.72 41.66 66.68 59.80 44.52
50% Palu chang2024palu 8.40 18.93 1.31 61.50 84.56 38.40 50.90 46.80 38.85
ReCalKV 8.39 18.89 1.37 58.50 84.75 39.41 58.29 54.61 40.53
60% Palu chang2024palu 5.10 16.51 2.13 55.50 59.84 33.13 29.62 33.56 29.42
ReCalKV 6.62 17.96 0.17 58.00 80.41 38.13 49.05 44.43 36.85
70% Palu chang2024palu 4.54 9.99 1.40 39.00 16.98 19.18 1.75 7.52 13.26
ReCalKV 3.28 15.41 0.12 53.00 66.24 32.61 34.11 32.15 29.62
LLaMA-2-13B-Chat llama2
0% Original 24.21 20.38 25.70 67.50 86.90 42.19 50.06 50.55 45.94
50% Palu chang2024palu 24.65 21.03 24.21 67.00 83.75 40.73 37.81 38.35 42.19
ReCalKV 19.30 20.47 24.32 68.00 83.82 40.96 29.72 36.41 40.38
60% Palu chang2024palu 17.65 20.27 21.76 65.00 79.25 36.49 34.04 29.91 38.05
ReCalKV 17.16 20.12 24.30 65.00 80.77 39.22 39.33 37.46 40.42
70% Palu chang2024palu 17.99 19.10 17.11 59.5 62.44 29.45 8.73 18.09 29.05
ReCalKV 16.47 20.22 22.18 62.50 76.87 35.50 30.79 26.51 36.38
Mistral-7B-Instruct-v0.2 mistral
0% Original 32.51 24.29 26.96 71.00 86.23 42.95 55.89 54.12 49.24
50% Palu chang2024palu 31.16 23.76 25.82 69.50 83.12 39.15 42.01 45.18 44.96
ReCalKV 31.71 23.35 26.43 70.00 82.51 39.22 46.12 45.21 45.57
60% Palu chang2024palu 21.21 23.73 24.75 68.00 76.59 36.14 26.24 30.48 38.39
ReCalKV 24.98 24.26 25.32 71.00 75.53 37.42 34.28 36.79 41.19
70% Palu chang2024palu 6.59 21.35 17.84 61.00 44.73 28.06 15.05 21.87 27.06
ReCalKV 9.13 23.01 20.85 65.00 51.44 31.37 15.13 22.66 29.82
LongChat-7B-v1.5-32k li2023long
0% Original 29.32 22.81 26.61 66.50 83.99 40.83 53.02 56.94 47.50
50% Palu chang2024palu 21.77 21.93 23.65 64.00 76.68 39.46 38.49 43.57 41.19
ReCalKV 25.15 22.08 23.38 63.00 79.75 40.72 50.54 50.52 44.39
60% Palu chang2024palu 13.12 21.97 19.07 55.50 66.14 34.68 42.01 16.55 33.63
ReCalKV 20.99 21.13 22.68 59.00 76.12 38.78 40.45 40.91 40.01
70% Palu chang2024palu 6.27 19.05 14.47 37.50 36.75 21.95 2.09 5.45 17.94
ReCalKV 17.50 20.70 18.94 44.00 67.29 33.86 10.48 15.23 28.50
4.2 Main Results
Perplexity Results. We evaluate the language modeling capability of ReCalKV on three standard datasets: WikiText2 wikitext2 , PTB plotz2017benchmarking , and C4 c4 , using perplexity as the evaluation metric. As shown in Table 1, ReCalKV achieves lower perplexity than Palu chang2024palu on most compression ratios and model families. On LLaMA-2-7B llama2 , ReCalKV yields a perplexity of 5.83 on WikiText2 and 8.14 on C4 at 50% compression, compared to 6.02 and 8.72 from Palu, respectively. Similar trends are observed on LLaMA-7B llama1 and Mistral-7B mistral . Notably, on PTB, ReCalKV significantly outperforms Palu under aggressive compression. For instance, at 70% compression, the perplexity of Palu rises sharply to 211.33 on LLaMA-7B and 172.23 on LongChat-7B, while ReCalKV keeps it much lower at 71.47 and 109.38, respectively. These results demonstrate that ReCalKV maintains stronger language modeling ability under high compression. Even at 70% compression, perplexity degradation remains moderate, indicating better information preservation than existing low-rank methods.
Zero-shot Accuracy Results. In addition to perplexity, we evaluate ReCalKV on six zero-shot QA datasets, including OBQA openbookqa , HellaSwag hellaswag , PIQA bisk2020piqa , ARC-e arc , ARC-c arc , and Winogrande winogrande . Across all model families and compression levels, ReCalKV demonstrates strong resilience in accuracy. While both methods see a decline in performance as the compression ratio increases, Palu chang2024palu exhibits a significantly steeper drop. For instance, at 70% compression on LLaMA-2-7B llama2 , Palu’s average accuracy drops to 52.14%, whereas ReCalKV retains 59.90%. Similar robustness is observed on Mistral-7B mistral and LongChat-7B li2023long , where ReCalKV consistently delivers higher or comparable average accuracy under the same compression levels. These results highlight ReCalKV’s strong capability to preserve task performance even under aggressive KV cache size reductions. Moreover, the method’s consistent stability across a wide range of tasks further underscores its practical utility for efficient long-context inference in real-world applications.
Longbench Result. We further evaluate ReCalKV on LongBench LongBench , a benchmark designed to test long-context understanding across diverse tasks. As shown in Table 2, ReCalKV achieves higher average accuracy than Palu chang2024palu across nearly all model scales and compression ratios. The gap becomes especially pronounced at high compression levels (e.g., 70%), where Palu suffers significant degradation while ReCalKV maintains competitive performance. This demonstrates ReCalKV’s robustness and suitability for long-sequence inference under memory constraints.
4.3 Ablation study
To analyze the individual contributions of each component in ReCalKV, we conduct ablation studies on LLaMA-2-7B llama2 under a fixed 80% compression ratio. Table 3 reports perplexity on WikiText-2 wikitext2 , PTB plotz2017benchmarking , and C4 c4 , as well as accuracy on two downstream evaluation suites: the average accuracy over six zero-shot QA datasets (zero-shot Avg. Acc) and the average accuracy across eight tasks from the LongBench benchmark LongBench (LongBench Avg. Acc).
Ablation on Head-wise Similarity-aware Reordering (HSR). By comparing the first and second rows in Table 3, we observe that enabling HSR alone (without offline calibration) significantly improves performance. For example, perplexity on WikiText-2 drops from 9.34 to 9.01, and LongBench accuracy increases from 9.01% to 12.44%. These results suggest that the reordering strategy in HSR effectively groups similar attention heads together before applying SVD, which reduces approximation error during low-rank decomposition and leads to improved model performance.
Ablation on Offline Calibration. Comparing the first and third rows, we assess the effect of offline calibration alone. Perplexity on WikiText-2 improves from 9.34 to 8.91, while LongBench accuracy rises to 13.09%. This confirms that calibrating the SVD decomposition of the Value projection matrix using a small held-out dataset effectively improves the quality of approximation, thereby enhancing model performance and robustness across tasks.
Table 3: Ablation studies on LLaMA-2-7B are conducted at a fixed 80% compression ratio. Perplexity is reported on WikiText-2, PTB, and C4, with our results in bold.
[…]
9.34 92.52 14.58 49.01 9.01
9.01 87.58 14.16 52.33 12.44
8.91 81.96 14.08 52.98 13.09
8.48 79.04 13.29 54.55 15.40
4.4 Integrate with KV Cache Quantization
Since ReCalKV reduces the hidden dimension of the KV cache, it is inherently orthogonal and thus complementary to KV cache quantization techniques, which compress along the bitwidth axis.
Table 4: ReCalKV with KV cache quantization.
Ratio Method Bit WikiText-2$\downarrow$ C4$\downarrow$
0% Original 16 5.47 7.26
50% Palu chang2024palu 4 6.04 8.75
Palu chang2024palu 3 6.15 8.92
ReCalKV 4 5.86 8.18
ReCalKV 3 5.96 8.34
60% Palu chang2024palu 4 6.84 10.77
Palu chang2024palu 3 7.01 11.06
ReCalKV 4 6.24 9.01
ReCalKV 3 6.39 9.21
70% Palu chang2024palu 4 8.71 15.17
Palu chang2024palu 3 9.04 15.75
ReCalKV 4 6.79 10.11
ReCalKV 3 7.01 10.41
This orthogonality enables seamless integration of the two methods for further compression. To assess their compatibility, we conduct experiments under various compression settings by combining ReCalKV with different quantization bitwidths (e.g., 4-bit and 3-bit), and varying the average rank to simulate different compression ratios. In our implementation, we adopt a per-token quantization scheme, allowing finer-grained control over dynamic range across sequence positions. Additionally, following Palu chang2024palu , we apply a randomized Hadamard transform prior to quantization to enhance robustness and reduce activation variance. Experimental results, as shown in Table 4, demonstrate that ReCalKV consistently outperforms Palu under the same compression ratio and bitwidth settings. For example, at a 60% compression ratio with 4-bit quantization, ReCalKV achieves a perplexity of 6.24 on WikiText-2, outperforming Palu’s 6.84. At 70% compression with 3-bit quantization, ReCalKV improves perplexity on C4 from 15.75 (Palu) to 10.41. These results confirm the effectiveness of combining ReCalKV with quantization for efficient KV cache compression.
5 Discussion
Limitations. While ReCalKV demonstrates high compression ratio with minimal accuracy drop, there remain several limitations to be addressed. First, despite our efforts to preserve performance, the compressed models still exhibit a certain degree of degradation in metrics such as perplexity and accuracy compared to the full KV cache baseline. This indicates that our method, like most lossy compression approaches, cannot fully maintain the original model behavior under high compression ratios. Second, ReCalKV currently focuses on hidden dimension reduction and does not incorporate token eviction strategies, which aim to reduce the KV cache length dimension. As these two directions are complementary, future work could explore the integration of ReCalKV with token eviction methods to further improve memory efficiency while maintaining performance.
Social Impacts. ReCalKV offers a practical and effective solution for compressing the KV cache in LLMs, significantly reducing memory consumption during long-context inference. This reduction in computational and memory requirements can lower both the deployment cost and energy consumption, thereby enabling wider access to LLMs in resource-constrained environments such as edge devices and developing regions. Although compression may lead to slight performance degradation, the trade-off is often acceptable in exchange for improved efficiency and scalability. By facilitating more sustainable and accessible AI deployment, ReCalKV contributes to the broader goal of inclusive and environmentally responsible AI development.
6 Conclusion
In this work, we prpose ReCalKV, a post-training KV cache compression framework tailored for efficient long-context reasoning in LLMs. By exploiting the distinct characteristics of Keys and Values in the attention mechanism, ReCalKV applies Head-wise Similarity-aware Reordering (HSR) and grouped SVD to compress Keys, while employing Offline Calibration and Matrix Fusion (OCMF) to compress Values. This design reduces hidden dimensions with minimal additional computation and preserves model performance under high compression ratios. Experimental results demonstrate that ReCalKV consistently outperforms existing low-rank compression methods, offering a practical and effective solution for memory-efficient LLM serving. Moreover, it can be combined with quantization to achieve higher compression with minimal performance loss. We believe this work provides a promising direction for scalable, long-context LLM deployment.